# A low-dimensional model for adaptive networks of spiking neurons

### v1

@@ Abstract
We analyze a large ensemble of Quadratic Integrate-and-Fire (QIF) neurons with heterogeneous input currents and adaptation variables. For a particular class of adaptation, which we call quadratic spike-frequency adaptation (QSFA), the high-dimensional system admits an exact reduction to a low-dimensional system of ordinary differential equations for three mean-field variables: the population’s firing rate, the mean membrane potential, and a mean adaptation variable. These low-dimensional firing rate equations (FRE) reveal an important generic feature of heterogeneous networks with spike frequency adaptation: Both the center and the width of the distribution of the neurons’ firing frequencies are reduced, and this largely favors the emergence of collective synchronization in the network. These results are confirmed by the bifurcation analysis of the FRE, which faithfully describe the collective dynamics of the spiking neuron network, including collective oscillations, bursting, and macroscopic chaos.
@@ I Introduction
However, important features of the microscopic dynamics related to the neuron-specific nature of SFA, which have significant implications at the population level, are not well described by such firing rate models. One overlooked effect, to the best of our knowledge, occurs in populations of neurons with SFA and heterogeneous firing frequencies (as in the models analyzed in Gast et al. (2020); Ferrara et al. (2023)): neurons with intrinsically high firing rates experience a greater reduction in firing frequency due to SFA than those with lower firing rates. Consequently, the overall level of frequency heterogeneity decreases, greatly favoring the emergence of collective synchronization Winfree (1967); Kuramoto (1984).
@@ II Populations of heterogeneous QIF neurons with spike-frequency adaptation (SFA)
The last two terms on the r.h.s. of Eq. (1a) vary from neuron to neuron and represent, respectively, constant inputs and adaptation currents (of strength $\beta\geq 0$). The definition of the QIF model requires a resetting rule such that after each spike —which is marked by the spike time $t_{j}^{k}$ at which $V_{j}$ reaches infinity—, the voltage is instantaneously reset to minus infinity. For the spike resetting at infinity, the spike-frequency (or firing rate) of intrinsically active neurons ($I_{j}-a_{j}>0$) is Ermentrout and Kopell (1986); Izhikevich (2007)
@@ II.2 Effect of QSFA on the distribution of firing frequencies
In Fig. 2, we show how this distribution changes with increasing adaptation strength $\beta>0$ for a Lorentzian distribution of inputs $I_{j}$, $g(I)=\Delta/\pi[(I-\bar{I})^{2}+\Delta^{2}]^{-1}$, of width $\Delta$, and centered at positive ($\bar{I}=2.5$) and negative ($\bar{I}=-1$) values. Increasing $\beta$ shifts the center of the distribution to the left, and reduces its width. Integration of $P_{0}(\nu)$ shows that the area under the graphs is independent of $\beta$. This indicates that QSFA does not alter the proportion of intrinsically spiking neurons in the population, which is solely determined by $g(I)$.
@@ III Firing Rate Equations With Quadratic Spike Frequency Adaptation
where $\eta_{j}$, $c_{j}$, and $\alpha_{j}(t)$ are all distributed according to Lorentzian probability density functions. Therefore, the model Eqs. (13) belongs to the class of models originally described in Montbrió et al. (2015), that are amenable to an exact, low-dimensional description in terms of two mean-field quantities, the population mean firing rate and membrane potential.
Next, we follow Montbrió et al. (2015) to derive a system of FRE that, in the thermodynamic limit $N\to\infty$, exactly describes the collective dynamics of the model Eqs. (13). Accordingly, we adopt the thermodynamic limit of Eqs. (1) and drop the indices in Eqs. (13,8). We denote by $\rho(V|\eta,c,t)$ the density of neurons with voltage $V$, given parameters $\eta$ and $c$, whose evolution is governed by the continuity equation
@@
Eqs. (17) and (18) couple all the Eqs. (16) with each other. By considering the analytic continuation of $W$ in the complex $\eta$ and $c$ planes, we require $\mathrm{Re}(W)$ to not become negative. We thus consider the poles of $g(\eta)$ and $f(c)$ such that $\partial_{t}\mathrm{Re}(W)|_{X=0}>0$, i.e. $\eta=\bar{\eta}-i\Delta$ and $c=\bar{c}+i\gamma$ 555 This can be seen by substituting Eq. 12 in Eq. 16 with $X(t)=R(t)=0$, which yields $\tau_{m}\partial_{t}W|_{X=0}=i\left[\eta-\eta h(t)-l(t)-ce^{-t/\tau}+Y^{2}\right].$ Then, for $X=\mathrm{Re}(W)$, the derivative $\partial_{t}X(t)$ evaluated at $X(t)=0$ can be evaluated using complex-valued $\eta$ and $c$, as $\tau_{m}\partial_{t}X|_{X=0}=-\eta_{i}(1-h)+c_{i}e^{-t/\tau}\stackrel{{\scriptstyle!}}{{>}}0.$ where the subscript $i$ indicates the imaginary parts of $\eta$ and $c$. Given that $h\in[0,1]$, this can be independently satisfied (either for $\eta_{i}\neq 0=c_{i}$, or for $c_{i}\neq 0=\eta_{i}$) only if $\eta_{i}<0$ or $c_{i}>0$. . Then, by applying Cauchy’s residue theorem, we find that
@@
with $A(0)=0$. After a transitory period of the order of $\tau$, the dynamics of Eq. 21 converges to the system of FRE
@@
The three-dimensional system Eqs. (24,23) governs the asymptotic collective dynamics of the population of QIF neurons Eqs. (1,2,6), where $A(t)$ corresponds to the mean of the adaptation variables $a_{j}(t)$ 666 After a transitory period of time $\tau$, the variables $A(t)$ and $|{B}(t)|$ correspond to the mean and the spread of the neurons’ adaptation variables $a_{j}(t)$. The solution Eq. 22 indicates that the width of the distribution of adaptation variables is reduced by the factor $\beta/(1+\beta)$. If $\bar{\eta}$ is time-independent, the solution of Eq. 23 can be written as $A(t)=\bar{\eta}(1-e^{-t\tau})/(1+\beta)+\tilde{A}(t)$ where $\tilde{A}(t)$ follows the dynamics $\tau_{a}\dot{\tilde{A}}=-\tilde{A}(1+\beta)+\beta J\tau_{m}R(t)$. Therefore, after a transient of the order $\tau$, the solution is $A(t)=\bar{\eta}/(1+\beta)+\tilde{A}(t)$, and the center of the distribution of adaptation variables is also reduced by a factor $\beta/(1+\beta)$. .
@@ IV Collective dynamics of populations of QIF neurons with QSFA
We next analyze the FREs (24,23) for globally coupled, excitatory QIF neurons with QSFA. We focus on the analysis of persistent states (PS), the onset of collective oscillations, as well as in the presence network bursts—that have been also found in spiking neuron networks with alternative models of SFA  Latham et al. (2000); Vreeswijk and Hansel (2001); Gigante et al. (2007); Dur-e Ahmad et al. (2012); Gast et al. (2020); Ferrara et al. (2023)—and collective chaos.
@@
where the population’s f-I curve is 777 Similar to the f-I curve of an individual QIF neuron with QSFA, Eq. (4), also the population f-I curve $\Phi$ is a non-negative, monotonously increasing function that scales as the square root of the input for large $I$.
@@
QSFA does not alter the shape of the f-I curve, but only scales it by a factor $1/\sqrt{1+\beta}$, which allows us to borrow the parametric formula for the SN boundaries in Montbrió et al. (2015) (corresponding to $\beta=0$) and use it for any value of $\beta$ 888 In parametric form, the SN curves —shown in Fig. 3— are: $\bar{\eta}_{SN}=-(1+\beta)(\pi\tau_{m}R^{*})^{2}-\frac{3\Delta^{2}}{4(1+\beta)(\pi\tau_{m}R^{*})^{2}},$ and $J_{SN}=2(1+\beta)\pi^{2}\tau_{m}R^{*}+\frac{\Delta^{2}}{2(1+\beta)\pi^{2}(\tau_{m}R^{*})^{3}}.$ The phase diagrams in Fig. 3 show two Saddle-Node (SN) bifurcation curves for various $\beta$ values, which meet in a cusp point. Within the region bounded by the SN bifurcations, asynchronous, low-activity states (LAS) coexist with PS. In the rest of the parameter space there exists a unique fixed point that represents an asynchronous state.
@@ V Conclusions
We demonstrate that this adaptation-induced homogenization significantly enhances the emergence of global oscillations in the network (see Fig. 2) compared to models without neuron-specific SFA. In those models, the synchronization region is confined to a narrower parameter space (cf. Fig. 2 in Ferrara et al. (2023)): The collective effect of SFA in models without neuron-specific adaptation variables reduces the intrinsic currents for each neuron by the same amount, turning intrinsically spiking neurons into quiescent neurons. This global reduction leads to the suppression of collective oscillations under strong adaptation conditions as the homogenization effect is lacking.
@@
Our numerical simulations of the exact FRE (24,23) closely follow those of the original network model Eqs. (1,2,6), as expected, see Fig. 6. Still, we advise caution when interpreting results from microscopic network simulations due to the presence of finite-size fluctuations. In the QSFA model (1b,2), the adaptation variable is allowed to take on negative values. This ‘negative adaptation’ increases the excitability of quiescent neurons by reducing the distance between their resting potential and the spiking threshold, allowing finite-size fluctuations to induce population bursts that would not occur in infinitely large networks, or in SFA models where adaptation is constrained to non-negative values.
Several studies have explored the mechanisms by which SFA synchronizes neural firing Vreeswijk and Hansel (2001); Fuhrmann et al. (2002); Ermentrout et al. (2001); Gigante et al. (2007). However, to our knowledge, the ability of SFA to reduce frequency heterogeneity within a neuronal population has been largely overlooked. Frequency heterogeneity plays a crucial role in determining the synchronization of a population of oscillators Winfree (1967); Kuramoto (1984). Therefore, we believe that the synchronization effects that we reported, which are described by the firing rate model Eqs. (24,23), may provide new insights into the mechanistic origins of neuronal oscillations.

### v2

# Low-dimensional model for adaptive networks of spiking neurons

@@ Abstract
We investigate a large ensemble of Quadratic Integrate-and-Fire (QIF) neurons with heterogeneous input currents and adaptation variables. Our analysis reveals that for a specific class of adaptation, termed quadratic spike-frequency adaptation (QSFA), the high-dimensional system can be exactly reduced to a low-dimensional system of ordinary differential equations, which describes the dynamics of three mean-field variables: the population’s firing rate, the mean membrane potential, and a mean adaptation variable. The resulting low-dimensional firing rate equations (FRE) uncover a key generic feature of heterogeneous networks with spike frequency adaptation: Both the center and the width of the distribution of the neurons’ firing frequencies are reduced, and this largely promotes the emergence of collective synchronization in the network. Our findings are further supported by the bifurcation analysis of the FRE, which accurately captures the collective dynamics of the spiking neuron network, including phenomena such as collective oscillations, bursting, and macroscopic chaos.
@@ I Introduction
However, key aspects of the microscopic dynamics associated with the neuron-specific nature of SFA are not adequately captured by such firing rate models. One overlooked phenomenon arises in populations of neurons with SFA and heterogeneous firing frequencies: neurons with intrinsically high firing rates undergo a more pronounced reduction in firing frequency due to SFA compared to neurons with lower firing rates. As a result, the overall level of frequency heterogeneity diminishes, significantly promoting the emergence of collective synchronization Winfree (1967); Kuramoto (1984).
@@ II Populations of heterogeneous QIF neurons with spike-frequency adaptation (SFA)
The last two terms on the r.h.s. of Eq. (1a) vary from neuron to neuron and represent, respectively, constant inputs and adaptation currents of strength $\beta\geq 0$. The definition of the QIF model requires a resetting rule such that after each spike —which is marked by the spike time $t_{j}^{k}$ at which $V_{j}$ reaches infinity—, the voltage is instantaneously reset to minus infinity. For the spike resetting at infinity, the spike-frequency (or firing rate) of intrinsically active neurons ($I_{j}-a_{j}>0$) is Ermentrout and Kopell (1986); Izhikevich (2007)
@@ II.2 Effect of QSFA on the distribution of firing frequencies
[…]
@@ III Firing Rate Equations With Quadratic Spike Frequency Adaptation
where $\eta_{j}$, $c_{j}$, and $\alpha_{j}(t)$ are all distributed according to Lorentzian probability density functions.
Equation (13) belongs to a class of mean field models that in the limit $N\to\infty$ admit an exact, low-dimensional description in terms of the population mean firing rate and membrane potential —see Eq. (19) in Ref. Montbrió et al. (2015). In the following we derive such low-dimensional FRE using the procedure originally proposed in Montbrió et al. (2015). Accordingly, we adopt the thermodynamic limit of Eqs. (1) and drop the indices in Eqs. (8,13). We denote by $\rho(V|\eta,c,t)$ the density of neurons with voltage $V$, given parameters $\eta$ and $c$, whose evolution is governed by the continuity equation
@@
Eqs. (17) and (18) couple the infinite set of Eqs. (16). By considering the analytic continuation of $W$ in the complex $\eta$ and $c$ planes, we require $\mathrm{Re}(W)$ to not become negative. We thus consider the poles of $g(\eta)$ and $f(c)$ such that $\partial_{t}\mathrm{Re}(W)|_{X=0}>0$, i.e. $\eta=\bar{\eta}-i\Delta$ and $c=\bar{c}+i\gamma$ 666 This can be seen by substituting Eq. 12 in Eq. 16 with $X(t)=R(t)=0$, which yields $\tau_{m}\partial_{t}W|_{X=0}=i\left[\eta-\eta h(t)-l(t)-ce^{-t/\tau}+Y^{2}\right].$ Then, for $X=\mathrm{Re}(W)$, the derivative $\partial_{t}X(t)$ evaluated at $X(t)=0$ can be evaluated using complex-valued $\eta$ and $c$, as $\tau_{m}\partial_{t}X|_{X=0}=-\eta_{i}(1-h)+c_{i}e^{-t/\tau}\stackrel{{\scriptstyle!}}{{>}}0.$ where the subscript $i$ indicates the imaginary parts of $\eta$ and $c$. Given that $h\in[0,1]$, this can be independently satisfied (either for $\eta_{i}\neq 0=c_{i}$, or for $c_{i}\neq 0=\eta_{i}$) only if $\eta_{i}<0$ or $c_{i}>0$. . Then, by applying Cauchy’s residue theorem, we find that
@@
with $A(0)=0$. Then, after a transitory period of the order of $\tau$, the dynamics of Eq. 21 converges to the system of FRE
@@
The three-dimensional system Eqs. (24,23) governs the asymptotic collective dynamics of the population of QIF neurons Eqs. (1,2,6), where $A(t)$ corresponds to the mean of the adaptation variables $a_{j}(t)$ 777 After a transitory period of time $\tau$, the variables $A(t)$ and $|{B}(t)|$ correspond to the mean and the spread of the neurons’ adaptation variables $a_{j}(t)$. The solution Eq. 22 indicates that the width of the distribution of adaptation variables is reduced by the factor $\beta/(1+\beta)$. If $\bar{\eta}$ is time-independent, the solution of Eq. 23 can be written as $A(t)=\bar{\eta}(1-e^{-t\tau})/(1+\beta)+\tilde{A}(t)$ where $\tilde{A}(t)$ follows the dynamics $\tau_{a}\dot{\tilde{A}}=-\tilde{A}(1+\beta)+\beta J\tau_{m}R(t)$. Therefore, after a transient of the order $\tau$, the solution is $A(t)=\bar{\eta}/(1+\beta)+\tilde{A}(t)$, and the center of the distribution of adaptation variables is also reduced by a factor $\beta/(1+\beta)$. .
@@ IV Collective dynamics of populations of QIF neurons with QSFA
We next analyze the FREs (24,23) for globally coupled, excitatory QIF neurons with QSFA. We focus on the analysis of persistent states (PS), the onset of collective oscillations, as well as in the presence of network bursts —that have been also found in spiking neuron networks with alternative models of SFA Latham et al. (2000); Vreeswijk and Hansel (2001); Gigante et al. (2007); Dur-e Ahmad et al. (2012); Gast et al. (2020); Ferrara et al. (2023)— and collective chaos.
@@
where the population’s f-I curve is 888 Similar to the f-I curve of an individual QIF neuron with QSFA, Eq. (4), also the population f-I curve $\Phi$ is a non-negative, monotonously increasing function that scales as the square root of the input for large $I$.
@@
QSFA does not alter the shape of the f-I curve, but only scales it by a factor $1/\sqrt{1+\beta}$, which allows us to borrow the parametric formula for the SN boundaries in Montbrió et al. (2015) (corresponding to $\beta=0$) and use it for any value of $\beta$ 999 In parametric form, the SN curves —shown in Fig. 3— are: $\bar{\eta}_{SN}=-(1+\beta)(\pi\tau_{m}R^{*})^{2}-\frac{3\Delta^{2}}{4(1+\beta)(\pi\tau_{m}R^{*})^{2}},$ and $J_{SN}=2(1+\beta)\pi^{2}\tau_{m}R^{*}+\frac{\Delta^{2}}{2(1+\beta)\pi^{2}(\tau_{m}R^{*})^{3}}.$ The phase diagrams in Fig. 3 show two Saddle-Node (SN) bifurcation curves for various $\beta$ values, which meet in a cusp point. Within the region bounded by the SN bifurcations, asynchronous, low-activity states (LAS) coexist with PS. In the rest of the parameter space there exists a unique fixed point that represents an asynchronous state.
@@ V Conclusions
Numerous studies have investigated the mechanisms by which SFA synchronizes neural firing Crook et al. (1998); Vreeswijk and Hansel (2001); Fuhrmann et al. (2002); Ermentrout et al. (2001); Gigante et al. (2007); Ladenbauer et al. (2012); Augustin et al. (2017). However, to the best of our knowledge, the potential of SFA to reduce frequency heterogeneity within a neuronal population has not been addressed. In Fig. Fig. 2, we show how this adaptation-induced homogenization markedly enhances the emergence of global oscillations in the network. We further confirm that the same homogenization effect occurs in networks of heterogeneous neurons with linear SFA, suggesting that synchronization is likely to be enhanced in these networks as well.
Additionally, we note that in mean field models of heterogeneous QIF neurons without neuron-specific SFA, the synchronization region is restricted to a narrower parameter space (cf. Fig. 2 in Ferrara et al. (2023)). In these models, SFA uniformly reduces the intrinsic currents across neurons, rendering intrinsically spiking neurons quiescent under strong adaptation conditions. This suppression of firing ultimately leads to the disappearance of synchronization in regimes of strong adaptation.
@@
Our numerical simulations of the exact FRE (24,23) closely follow those of the original network model Eqs. (1,2,6), as expected, see Fig. 5. Still, we advise caution when interpreting results from microscopic network simulations due to the presence of finite-size fluctuations. In the QSFA model (1b,2), the adaptation variable is allowed to take on negative values. This ‘negative adaptation’ increases the excitability of quiescent neurons by reducing the distance between their resting potential and the spiking threshold, allowing finite-size fluctuations to induce population bursts that would not occur in infinitely large networks, or in SFA models where adaptation is constrained to non-negative values
Finally, an interesting direction for future research would be to consider populations of neurons subject to stochastic inputs. Recent studies have investigated how different types of noise can be incorporated and analyzed within the theoretical framework of mean field models, such as the one investigated here, see e.g. Clusella and Montbrió (2024); Pietras et al. (2023); Pyragas and Pyragas (2023); Goldobin et al. (2024, 2021); Goldobin (2021); Di Volo et al. (2022); Klinshov and Kirillov (2022); Goldobin et al. (2023); Vinci et al. (2023). However, it remains an open question whether some of these findings can be extended to the QIF model with QSFA proposed here.